OpenStudy (anonymous):

find the center and radius of the circle and identify the domain and range. x^2 + y^2 +2y=3

OpenStudy (anonymous):

start with \[x^2+(y+1)^2=3+1=4\] and then read off the solution

OpenStudy (anonymous):

Put the equation into standard form by completing the square.\[x^2+(y+1)^2=4\]Are the answers a little more obvious now?

OpenStudy (anonymous):

yes that makes more sense :)

OpenStudy (anonymous):

yay!